# 70% Size, 100% Accuracy: Lossless LLM Compression for Efficient GPU Inference via Dynamic-Length Float

### v2

@@ 1 Introduction
Model compression via quantization has limitations. Quantization is a type of lossy compression method that lowers the precision of model weights by converting them into lower bit-width representations frantar2022gptq ; lin2024awq ; li2023q ; sui2024bitsfusion . Although it can significantly reduce memory usage and often improve inference speed, quantization is not a one-size-fits-all solution and presents several key limitations: ➊ Accuracy degradation. By design, quantization introduces approximation errors. The degree of accuracy loss depends on multiple factors, including the base model, quantization method, evaluation benchmark, and target bit-width eval_quantized . These interactions make it difficult to predict or quantify the impact comprehensively. Even mild quantization can noticeably degrade performance. For example, applying 8-bit SmoothQuant xiao2023smoothquant to DeepSeek-R1-Distill-Qwen-1.5B guo2025deepseek results in a 9.09% drop in average accuracy across reasoning tasks liu2025quant_reason_bench . ➋ Behavioral shifts. Even when overall accuracy metrics appear roughly unchanged, quantized models may behave differently from their full-precision counterparts. For instance, Dutta et al. dutta2024accuracy_not observe a phenomenon called flips, where quantized models produce answers that change from correct to incorrect and vice versa. This indicates that quantization can significantly alter model behavior, even when standard accuracy metrics show minimal change. For example, the W8A16 GPTQ-quantized Qwen2-1.5Bfrantar2022gptq ; yang2024qwen2 exhibits only a 0.3% drop in GSM8K (8-shot) accuracy cobbe2021gsm8k , yet 6.37% of its answers flip in correctness dutta2024accuracy_not . ➌ Compliance and reliability concerns. In domains like finance or healthcare, quantized models may not satisfy regulatory or reliability standards, as their outputs may differ from those of the original models kharinaev2025invest_quant_safety . We refer readers to Appendix A for a more detailed discussion on quantization.
@@
Existing lossless model compression does not support efficient GPU inference. Unlike lossy compression, lossless compression reduces model size while preserving the full precision of the original weights. This ensures the model’s output distribution remains identical to that of the uncompressed counterpart. However, most existing lossless methods focus on storage efficiency, such as compressing model checkpoints han2015deep_comp ; hershcovitch2024zipnn , or target specialized hardware like FPGAs yubeaton2025huffllm , rather than accelerating inference on general-purpose GPUs. While useful for tasks like checkpoint rollback during large-scale training wang2023gemini_ckpt or reducing download time from model hubs hershcovitch2024zipnn , these methods offer little to no benefit for GPU-based inference.
@@
Our proposal, Dynamic-Length Float (DFloat11), is a lossless compression framework optimized for efficient GPU inference. We identify a key inefficiency in the commonly used BFloat16 format: its 8-bit exponent field carries only about 2.6 bits of actual information. This redundancy is consistent across a wide range of LLMs, as shown in Section 2.2. To exploit it, we apply Huffman coding huffman_coding to the exponent bits of BFloat16 weights, while leaving the sign and mantissa bits uncompressed. The resulting exponents have dynamic-length encodings: frequent values are assigned shorter codes, while rarer ones use longer codes. However, standard Huffman decoding relies on sequential bit-by-bit tree traversal, which is inefficient on GPUs due to limited parallelism. Assigning one GPU thread per decompression task leads to severe hardware underutilization and high latency. To overcome this, we design a hardware-aware algorithm that enables efficient online decompression of dynamic-length floats on GPUs. Our solution includes three key components: 1. compact, hierarchical lookup tables (LUTs) that fit in GPU SRAM to support fast, table-based Huffman decoding, 2. a two-phase GPU kernel that uses lightweight auxiliary variables to coordinate thread-level read and write operations, and 3. batched decompression at the transformer-block level to maximize throughput. We summarize our contributions as follows:
@@ 2.1 Preliminary
Brain Float (BFloat16) Recent state-of-the-art LLMs predominantly employ the 16-bit Brain Float format (BFloat16 or BF16) for storing weights, due to its balance of numerical precision with memory efficiency. BF16 allocates its 16 bits as follows: 1 sign bit, 8 exponent bits, and 7 mantissa bits. The numerical value represented by a BF16 number is computed as:
@@
Entropy Coding Entropy coding is a core technique in lossless data compression that leverages statistical redundancy to reduce data size. Several widely used methods fall under this category, including Huffman coding huffman_coding , arithmetic coding arithmetic_coding , and Asymmetric Numeral Systems (ANS) ans . Among these, Huffman coding is one of the most widely adopted, which uses variable-length encoding to minimize the size of encoded data. It assigns shorter binary codes to more frequent symbols and longer codes to less frequent ones. The codes are decoded using a prefix-free binary tree, known as a Huffman tree. Due to the prefix-free property of Huffman codes, no code is a prefix of any other, which ensures unique decodability of the encoded bitstream without the need for delimiters. The tree is constructed based on symbol frequencies and is provably optimal for any given frequency distribution. However, decoding Huffman codes in a massively parallel manner is challenging due to its inherently sequential nature.
GPU Computation and Memory Paradigm GPUs are designed to perform computations in a massively parallel manner. A modern GPU consists of thousands of threads, which are organized into blocks and executed on streaming multiprocessors (SMs). Each block has access to a small, fast, on-chip memory called shared memory (often referred to as SRAM), which provides much lower latency and higher bandwidth than the off-chip global memory, commonly known as high-bandwidth memory (HBM). The capacity of shared memory is limited, typically having up to 100 KB per block. In this work, we leverage the fast access characteristics of SRAM to enable efficient on-the-fly decompression of compressed weights during inference.
@@ 2.3 Dynamic-Length Float: Lossless LLM Compression for Efficient GPU Inference
The Core Challenge: Efficient GPU Inference with Compressed Weights While DFloat11 enables lossless compression of LLM weights, efficient GPU inference remains a key challenge. Entropy-coded weights use variable-length encoding and cannot be directly used in matrix multiplications. As a result, each weight matrix must be decompressed on-the-fly to its original BFloat16 format before matrix multiplication, then discarded immediately after use to conserve memory. However, traditional Huffman decoding is inherently sequential, requiring bit-by-bit tree traversal for each element, which is ill-suited for GPUs’ parallel architecture. Naively assigning a single thread for decompression leads to poor utilization and high latency. Addressing this bottleneck is essential for practical compressed inference.
@@ 2.3.1 Efficient Decoding with Hierarchical Lookup Tables
The traditional approach to decoding Huffman codes involves reading the encoded bitstream bit by bit and traversing the Huffman tree accordingly. However, this method is inefficient on GPUs due to frequent branching and thread divergence. To enable efficient decoding on GPUs, we adopt a lookup-table-based approach huffman_gap_arrays .
@@ 3.1 Results
Figure 4: Comparison of throughput (left three) and latency (right three) for token decoding using the original BF16 models and their DF11-compressed counterparts. Portions of the BF16 models are offloaded to the CPU due to GPU memory constraints.
@@
DF11 compresses models to 70% size. Table 1 presents the compression factors of DF11 for a wide selection of recent LLMs and DMs. Specifically, we apply compression to all weight matrices and token embeddings in LLMs and all weight matrices in the transformer blocks of DMs. The models we compress include Llama 3.1/3.3 grattafiori2024llama , Qwen 3 yang2024qwen2 , Mistral Nemo/Small mistral2024nemo ; mistral2025small3 , Phi 4 phi4 , DeepSeek R1 Distilled guo2025deepseek , Stable Diffusion 3.5 stablediffusion35 , FLUX.1 flux . DF11 achieves approximately 70% compression across all models, corresponding to an effective bit width of around 11 bits.
Accuracy and perplexity evaluations confirm DF11 compression is lossless. We verify the lossless property of DF11 compression through a series of accuracy and perplexity evaluations on standard benchmarks. Evaluations are conducted using lm_evaluation_harness eval-harness , reporting accuracy on MMLU mmlu and TruthfulQA truthfulqa , and word-level perplexity on WikiText wikitext and C4 c4 . The results are shown in Table 2. As demonstrated, the compressed model achieves identical accuracy and perplexity to the original BF16 counterpart. We also present the text-to-image generation results of BF16 and DF11 Stable Diffusion 3.5 Large model in Appendix J. Given the same random seed and text prompt, the image generated are pixel-wise identical with the original model.
@@
DF11 reduces memory usage for diffusion transformers with minimal latency impact. We evaluate the effect of DF11 compression on diffusion transformer models by measuring peak GPU memory and text-to-image generation time for an $\times1024$ image, over 5 runs. As shown in Table 3, DF11 reduces memory consumption by 28.3% for Stable Diffusion 3.5 and 27.8% for FLUX.1. The relative increase in latency is small: 4.1% for Stable Diffusion and 5.5% for FLUX.1.
@@ 5 Conclusion
We present Dynamic-Length Float (DFloat11), a lossless compression framework for efficient GPU inference of BFloat16 models. DFloat11 leverages the information inefficiency of foundation model weights through entropy-coded, dynamic-length encoding, achieving compression near the information-theoretic limit. We develop hardware-aware algorithms to support fast inference with compressed weights. Extensive evaluations show that DFloat11 lowers GPU memory requirement for efficiently running LLMs and DMs, enables longer generation lengths, while introducing minimal decompression overhead and preserving bit-exact accuracy.
@@ Appendix A Discussion: Is Quantization a Universal Solution?
Much of the motivation behind our work lies in understanding whether lossless compression of large-scale models such as LLMs, which preserves 100% identical output behavior compared to the original uncompressed model, is a practical direction worthy of further study. Specifically, how does DFloat11, which compresses LLMs to approximately 11 bits, compare to widely used lossy quantization techniques [14, 32], where models are typically reduced to even lower bit-widths (e.g., 8-bit or 4-bit)?
The answer is far more nuanced than a simple “Yes/No” or a one-size-fits-all judgment about which approach is better. For instance, existing benchmark studies like [17, 46, 25] often suggest that 8-bit (weight-only or not) quantization is a relatively “safe” compression scheme. Although technically lossy, 8-bit models can often maintain strong task performance across a range of standard benchmarks. However, we must note these benchmarks typically focus on a narrow set of tasks (e.g., WikiText2 perplexity, MMLU, Commonsense Reasoning), and thus fail to offer a comprehensive view of real-world LLM usage, especially from the perspective of end-users.
That being said, the argument that “current benchmarks fail to capture the performance gap between 8-bit compressed and 16-bit uncompressed models” is itself constrained by the limitations of the current benchmarking landscape, making it difficult to produce abundant supporting evidence. Nonetheless, some reports have begun to highlight such gaps. For example, human evaluations on LLM Arena [^1] https://x.com/lmarena_ai/status/1835760196758728898 show a notable performance drop between Llama-3.1-405B-Instruct [18] and its W8A8 counterpart (Llama-3.1-405B-Instruct-FP8), particularly under coding (1293 vs. 1277) and long-query (1282 vs. 1275) tasks. Similarly, quantizing DeepSeek-R1-Distill-Llama-70B [19] from 16 bits to 8 bits results in a 23.7% drop on GPQA (from 9.51% to 7.25%). [^2] https://huggingface.co/RedHatAI/DeepSeek-R1-Distill-Llama-70B-quantized.w8a8 Furthermore, reasoning, a core capability of modern LLMs, appears especially sensitive to compression loss. Recent benchmark [34] reveals that quantizing DeepSeek-R1-Distill-Qwen-1.5B with 8-bit SmoothQuant [43] (for weight, attention, and KV cache) leads to an average 9.09% drop in reasoning tasks (48.82% to 44.29%) across datasets like AIME, MATH-500, GPQA-Diamond, and LiveCodeBench. We leave more evidence exploring the performance gap between 8-bit quantized and uncompressed model in Appendix H.
@@
To eliminate this burden, DFloat11 offers a compelling alternative: delivering 100% identical performance to the original model, while consuming only $\sim$70% of the memory footprint with many throughput benefits, which is a unique and practical offering for resource-constrained deployment settings.
@@ Appendix B Extended Related Works
LLM weights are typically stored in compact floating-point formats such as FP16 or BF16. FP16 allocates 1 sign bit, 5 exponent bits, and 10 mantissa bits, whereas BF16 uses 1 sign bit, 8 exponent bits, and 7 mantissa bits. Compared to FP16, BF16 offers a wider dynamic range at the cost of precision, which improves numerical stability and mitigates overflow issues during training [15, 26].
Compressed data formats typically aim for lower bit-widths. For example, FP8—which comes in both E4M3 (4 exponent bits, 3 mantissa bits, plus 1 sign bit) and E5M2 configurations—has seen reasonable adoption in LLM training and development. Integer formats like INT8 have also been well explored, as in LLM.int8() [7] and its following works. Formats with a stronger emphasis on efficiency, such as FP4, INT4, NF4 [8], and AF4 [49], use only 4 bits. In this work, we primarily focus on formats with $\geq$8 bits, as benchmark literature [46, 17, 34] often suggests that 8-bit quantization results in negligible performance drop—though we show in Section A that this claim is likely skewed due to evaluation selectiveness and benchmark limitations.
@@
While lossy model compression techniques such as pruning and quantization [13, 32, 14] have received widespread attention, lossless model compression remains a relatively underexplored area. Upon careful investigation, we identified roughly four prior works that have made meaningful efforts in this space. Deep Compression [20] is a foundational work, applying Huffman coding [24] to quantized CNN models and achieving an additional $\sim$22% compression gain for model checkpoints. ZipNN [23] extended this idea to language models, comparing its results to classic lossless compression tools such as zlib [9] and zstd [^3] https://github.com/facebook/zstd and demonstrated superior compression gains. However, this line of work—including their industry counterparts, such as ezm7444https://github.com/liuliu/s4nnc/pull/11 https://encode.su/threads/4067-Good-Compressors-for-16-bit-floats—is limited in that its efficiency gains only apply to storage (reducing the size of model checkpoints) but offer no benefits during inference. While such storage savings are meaningful in large-scale training settings—where frequent snapshotting and checkpoint rollbacks are needed [41]—they have limited impact for everyday LLM end-users. Model downloading is typically a one-time cost, so even if a model checkpoint is compressed by 50%, it only cuts the download time at most by half, presumably over the model’s entire lifecycle of deployment. Furthermore, checkpoints are usually stored on disk, where terabytes of capacity are easily available, making up a much looser constraint compared to GPU HBM (High Bandwidth Memory); one of the main resource constraints during inference.
@@
We argue that a lossless compression technique would be substantially more impactful if it could deliver efficiency gains during inference—particularly on GPU-based systems, which is the default setup for LLM serving. In this context, NeuZip [21] is the only prior work we identify that supports GPU inference. NeuZip applies entropy encoding with layer-wise decompression to maintain a reduced memory footprint throughout serving. However, it is built on NVIDIA’s nvCOMP: “a high-speed data compression and decompression library optimized for NVIDIA GPUs”. [^5] https://developer.nvidia.com/nvcomp Unfortunately, nvCOMP is no longer open-source (only binary executables are available), which hinders future research. Moreover, we empirically find that nvCOMP’s inference throughput and latency are significantly worse than our proposed DFloat11 kernel, resulting in a pipeline that trades memory efficiency for substantial inference overhead (see Figure 7).
@@
Another work referencing NeuZip is Huff-LLM [50], which also aims to reduce memory costs while maintaining efficient inference. However, its contributions are specific to FPGA-like architectures and do not apply to GPUs. To the best of our knowledge, the DFloat data format we presented (and its respective kernel support in DFloat11) shall serve as the only GPU-inference-friendly data format with lossless compression benefits.

### v3

@@ 1 Introduction
Quantization is a type of lossy compression method that lowers the precision of model weights by converting them into lower bit-width representations frantar2022gptq; lin2024awq; li2023q; sui2024bitsfusion. Although it can significantly reduce memory usage and often improve inference speed, quantization is not a one-size-fits-all solution and presents several key limitations: ➊ Accuracy degradation. By design, quantization introduces approximation errors. The degree of accuracy loss depends on multiple factors, including the base model, quantization method, evaluation benchmark, and target bit-width eval_quantized. These interactions make it difficult to predict or quantify the impact comprehensively. Even mild quantization can noticeably degrade performance. For example, applying 8-bit SmoothQuant xiao2023smoothquant to DeepSeek-R1-Distill-Qwen-1.5B guo2025deepseek results in a 9.09% drop in average accuracy across reasoning tasks liu2025quant_reason_bench. ➋ Behavioral shifts. Even when overall accuracy metrics appear roughly unchanged, quantized models may behave differently from their full-precision counterparts. For instance, Dutta et al. dutta2024accuracy_not observe a phenomenon called flips, where quantized models produce answers that change from correct to incorrect and vice versa. This indicates that quantization can significantly alter model behavior, even when standard accuracy metrics show minimal change. For example, the W8A16 GPTQ-quantized Qwen2-1.5Bfrantar2022gptq; yang2024qwen2 exhibits only a 0.3% drop in GSM8K (8-shot) accuracy cobbe2021gsm8k, yet 6.37% of its answers flip in correctness dutta2024accuracy_not. ➌ Compliance and reliability concerns. In domains like finance or healthcare, quantized models may not satisfy regulatory or reliability standards, as their outputs may differ from those of the original models kharinaev2025invest_quant_safety. We refer readers to Appendix A for a more detailed discussion on quantization.
Existing lossless model compression does not support efficient GPU inference.
@@
Unlike lossy compression, lossless compression reduces model size while preserving the full precision of the original weights. This ensures the model’s output distribution remains identical to that of the uncompressed counterpart. However, most existing lossless methods focus on storage efficiency, such as compressing model checkpoints han2015deep_comp; hershcovitch2024zipnn, or target specialized hardware like FPGAs yubeaton2025huffllm, rather than accelerating inference on general-purpose GPUs. While useful for tasks like checkpoint rollback during large-scale training wang2023gemini_ckpt or reducing download time from model hubs hershcovitch2024zipnn, these methods offer little to no benefit for GPU-based inference.
Our proposal, Dynamic-Length Float (DFloat11), is a lossless compression framework optimized for efficient GPU inference.
@@
We identify a key inefficiency in the commonly used BFloat16 format: its 8-bit exponent field carries only about 2.6 bits of actual information. This redundancy is consistent across a wide range of LLMs, as shown in Section 2.2. To exploit it, we apply Huffman coding huffman_coding to the exponent bits of BFloat16 weights, while leaving the sign and mantissa bits uncompressed. The resulting exponents have dynamic-length encodings: frequent values are assigned shorter codes, while rarer ones use longer codes. However, standard Huffman decoding relies on sequential bit-by-bit tree traversal, which is inefficient on GPUs due to limited parallelism. Assigning one GPU thread per decompression task leads to severe hardware underutilization and high latency. To overcome this, we design a hardware-aware algorithm that enables efficient online decompression of dynamic-length floats on GPUs. Our solution includes three key components: 1. compact, hierarchical lookup tables (LUTs) that fit in GPU SRAM to support fast, table-based Huffman decoding, 2. a two-phase GPU kernel that uses lightweight auxiliary variables to coordinate thread-level read and write operations, and 3. batched decompression at the transformer-block level to maximize throughput. We summarize our contributions as follows:
@@ 2.1 Preliminary
Brain Float (BFloat16)
Recent state-of-the-art LLMs predominantly employ the 16-bit Brain Float format (BFloat16 or BF16) for storing weights, due to its balance of numerical precision with memory efficiency. BF16 allocates its 16 bits as follows: 1 sign bit, 8 exponent bits, and 7 mantissa bits. The numerical value represented by a BF16 number is computed as:
@@
Entropy Coding
Entropy coding is a core technique in lossless data compression that leverages statistical redundancy to reduce data size. Several widely used methods fall under this category, including Huffman coding huffman_coding, arithmetic coding arithmetic_coding, and Asymmetric Numeral Systems (ANS) ans. Among these, Huffman coding is one of the most widely adopted, which uses variable-length encoding to minimize the size of encoded data. It assigns shorter binary codes to more frequent symbols and longer codes to less frequent ones. The codes are decoded using a prefix-free binary tree, known as a Huffman tree. Due to the prefix-free property of Huffman codes, no code is a prefix of any other, which ensures unique decodability of the encoded bitstream without the need for delimiters. The tree is constructed based on symbol frequencies and is provably optimal for any given frequency distribution. However, decoding Huffman codes in a massively parallel manner is challenging due to its inherently sequential nature.
GPU Computation and Memory Paradigm
GPUs are designed to perform computations in a massively parallel manner. A modern GPU consists of thousands of threads, which are organized into blocks and executed on streaming multiprocessors (SMs). Each block has access to a small, fast, on-chip memory called shared memory (often referred to as SRAM), which provides much lower latency and higher bandwidth than the off-chip global memory, commonly known as high-bandwidth memory (HBM). The capacity of shared memory is limited, typically having up to 100 KB per block. In this work, we leverage the fast access characteristics of SRAM to enable efficient on-the-fly decompression of compressed weights during inference.
@@ 2.3 Dynamic-Length Float: Lossless LLM Compression for Efficient GPU Inference
The Core Challenge: Efficient GPU Inference with Compressed Weights
While DFloat11 enables lossless compression of LLM weights, efficient GPU inference remains a key challenge. Entropy-coded weights use variable-length encoding and cannot be directly used in matrix multiplications. As a result, each weight matrix must be decompressed on-the-fly to its original BFloat16 format before matrix multiplication, then discarded immediately after use to conserve memory. However, traditional Huffman decoding is inherently sequential, requiring bit-by-bit tree traversal for each element, which is ill-suited for GPUs’ parallel architecture. Naively assigning a single thread for decompression leads to poor utilization and high latency. Addressing this bottleneck is essential for practical compressed inference.
@@ 2.3.1 Efficient Decoding with Hierarchical Lookup Tables
The traditional approach to decoding Huffman codes involves reading the encoded bitstream bit by bit and traversing the Huffman tree accordingly. However, this method is inefficient on GPUs due to frequent branching and limited parallelism. To enable efficient decoding on GPUs, we adopt a lookup-table-based approach huffman_gap_arrays.
@@ 3.1 Results
Figure 4: Comparison of throughput (top row) and latency (bottom row) for token decoding using the original BF16 models and their DF11-compressed counterparts. Portions of the BF16 models are offloaded to the CPU due to GPU memory constraints.
@@
DF11 compresses models to 70% size.
Table 1 presents the compression factors of DF11 for a wide selection of recent LLMs and DMs. Specifically, we apply compression to all weight matrices and token embeddings in LLMs and all weight matrices in the transformer blocks of DMs. The models we compress include Llama 3.1/3.3 grattafiori2024llama, Qwen 3 yang2024qwen2, Mistral Nemo/Small mistral2024nemo; mistral2025small3, Phi 4 phi4, DeepSeek R1 Distilled guo2025deepseek, Stable Diffusion 3.5 stablediffusion35, FLUX.1 flux. DF11 achieves approximately 70% compression across all models, corresponding to an effective bit width of around 11 bits.
Accuracy and perplexity evaluations confirm DF11 compression is lossless.
We verify the lossless property of DF11 compression through a series of accuracy and perplexity evaluations on standard benchmarks. Evaluations are conducted using lm_evaluation_harness eval-harness, reporting accuracy on MMLU mmlu and TruthfulQA truthfulqa, and word-level perplexity on WikiText wikitext and C4 c4. The results are shown in Table 2. As demonstrated, the compressed model achieves identical accuracy and perplexity to the original BF16 counterpart. We also present the text-to-image generation results of BF16 and DF11 Stable Diffusion 3.5 Large model in Appendix J. Given the same random seed and text prompt, the image generated are pixel-wise identical with the original model.
DF11 outperforms CPU offloading in inference efficiency.
@@
DF11 reduces memory usage for diffusion transformers with minimal latency impact.
We assess the impact of DF11 compression on diffusion transformer models by measuring peak GPU memory usage and text-to-image generation latency for an $\times1024$ image across five runs. Neither the BF16 nor DF11 models employ CPU offloading. As shown in Table 3, DF11 reduces memory consumption by 28.3% for Stable Diffusion 3.5 and 27.8% for FLUX.1. The relative increase in latency is small: 4.1% for Stable Diffusion and 5.5% for FLUX.1.
DF11 memory savings enable longer generation lengths.
@@ 5 Conclusion
We introduce Dynamic-Length Float (DFloat11), a lossless compression framework designed for efficient GPU inference of BFloat16 models, including both large language models (LLMs) and diffusion models (DMs). DFloat11 exploits the information redundancy inherent in foundation model weights through entropy-coded, dynamic-length encoding, achieving compression rates close to the information-theoretic limit. To enable efficient deployment, we develop hardware-aware algorithms that support high-speed inference directly on compressed weights. Extensive experiments demonstrate that DFloat11 significantly reduces GPU memory requirements for LLMs and DMs, allowing for longer generation lengths, while maintaining bit-exact accuracy and incurring only negligible decompression overhead.
Acknowledgements
This work was supported by National Science Foundation SHF-2211815 and Ken Kennedy Institute Cluster Grants. Additionally, Henry and Xia are supported by ITE-2429680, IIS-2310260, and US Department of Transportation (USDOT) Tier-1 University Transportation Center (UTC) Transportation Cybersecurity Center for Advanced Research and Education (CYBER-CARE) grant #69A3552348332. Mohsen and Vipin are supported by OAC-2320952, OAC-2112606, and OAC-2117439. The views and conclusions in this paper are those of the authors and do not represent the views of any funding or supporting agencies.
@@ Appendix A Discussion: Is Quantization a Universal Solution?
Much of the motivation behind our work lies in understanding whether lossless compression of large-scale models such as LLMs, which preserves 100% identical output behavior compared to the original uncompressed model, is a practical direction worthy of further study. Specifically, how does DFloat11, which compresses LLMs to approximately 11 bits, compare to widely used lossy quantization techniques [frantar2022gptq, lin2024awq], where models are typically reduced to even lower bit-widths (e.g., 8-bit or 4-bit)?
The answer is far more nuanced than a simple “Yes/No” or a one-size-fits-all judgment about which approach is better. For instance, existing benchmark studies like [gong2024llmc, yang2024llmcbench, jin2024comp_eval_quant] often suggest that 8-bit (weight-only or not) quantization is a relatively “safe” compression scheme. Although technically lossy, 8-bit models can often maintain strong task performance across a range of standard benchmarks. However, we must note these benchmarks typically focus on a narrow set of tasks (e.g., WikiText2 perplexity, MMLU, Commonsense Reasoning), and thus fail to offer a comprehensive view of real-world LLM usage, especially from the perspective of end-users.
That being said, the argument that “current benchmarks fail to capture the performance gap between 8-bit compressed and 16-bit uncompressed models” is itself constrained by the limitations of the current benchmarking landscape, making it difficult to produce abundant supporting evidence. Nonetheless, some reports have begun to highlight such gaps. For example, human evaluations on LLM Arena [^1] https://x.com/lmarena_ai/status/1835760196758728898 show a notable performance drop between Llama-3.1-405B-Instruct [grattafiori2024llama] and its W8A8 counterpart (Llama-3.1-405B-Instruct-FP8), particularly under coding (1293 vs. 1277) and long-query (1282 vs. 1275) tasks. Similarly, quantizing DeepSeek-R1-Distill-Llama-70B [guo2025deepseek] from 16 bits to 8 bits results in a 23.7% drop on GPQA (from 9.51% to 7.25%). [^2] https://huggingface.co/RedHatAI/DeepSeek-R1-Distill-Llama-70B-quantized.w8a8 Furthermore, reasoning, a core capability of modern LLMs, appears especially sensitive to compression loss. Recent benchmark [liu2025quant_reason_bench] reveals that quantizing DeepSeek-R1-Distill-Qwen-1.5B with 8-bit SmoothQuant [xiao2023smoothquant] (for weight, attention, and KV cache) leads to an average 9.09% drop in reasoning tasks (48.82% to 44.29%) across datasets like AIME, MATH-500, GPQA-Diamond, and LiveCodeBench. We leave more evidence exploring the performance gap between 8-bit quantized and uncompressed model in Appendix H.
@@
To eliminate this burden, DFloat11 offers a compelling alternative: delivering 100% identical performance to the original model, while consuming only $\sim$70% of the memory footprint with throughput benefits, which is a unique and practical offering for resource-constrained deployment settings.
@@ Appendix B Extended Related Works
LLM weights are typically stored in compact floating-point formats such as FP16 or BFloat16 (officially stylized as bfloat16333https://cloud.google.com/blog/products/ai-machine-learning/bfloat16-the-secret-to-high-performance-on-cloud-tpus). FP16 allocates 1 sign bit, 5 exponent bits, and 10 mantissa bits, whereas BFloat16 uses 1 sign bit, 8 exponent bits, and 7 mantissa bits. Compared to FP16, BFloat16 offers a wider dynamic range at the cost of precision, which improves numerical stability and mitigates overflow issues during training [fujii2024fp8vsbf16, kalamkar2019studybf16].
Compressed data formats typically aim for lower bit-widths. For example, FP8—which comes in both E4M3 (4 exponent bits, 3 mantissa bits, plus 1 sign bit) and E5M2 configurations—has seen reasonable adoption in LLM training and development. Integer formats like INT8 have also been well explored, as in LLM.int8() [dettmers2022llmint8] and its following works. Formats with a stronger emphasis on efficiency, such as FP4, INT4, NF4 [dettmers2023qlora], and AF4 [yoshida2023af4], use only 4 bits. In this work, we primarily focus on formats with $\geq$8 bits, as benchmark literature [yang2024llmcbench, gong2024llmc, liu2025quant_reason_bench] often suggests that 8-bit quantization results in negligible performance drop—though we show in Section A that this claim is likely skewed due to evaluation selectiveness and benchmark limitations.
@@
While lossy model compression techniques such as pruning and quantization [frantar2023sparsegpt, lin2024awq, frantar2022gptq] have received widespread attention, lossless model compression remains a relatively underexplored area. Upon careful investigation, we identified roughly four prior works that have made meaningful efforts in this space. Deep Compression [han2015deep_comp] is a foundational work, applying Huffman coding [huffman_coding] to quantized CNN models and achieving an additional $\sim$22% compression gain for model checkpoints. ZipNN [hershcovitch2024zipnn] extended this idea to language models, comparing its results to classic lossless compression tools such as zlib [deutsch1996zlib] and zstd [^4] https://github.com/facebook/zstd and demonstrated superior compression gains. However, this line of work—including their industry counterparts, such as ezm7555https://github.com/liuliu/s4nnc/pull/11 https://encode.su/threads/4067-Good-Compressors-for-16-bit-floats—is limited in that its efficiency gains only apply to storage (reducing the size of model checkpoints) but offer no benefits during inference. While such storage savings are meaningful in large-scale training settings—where frequent snapshotting and checkpoint rollbacks are needed [wang2023gemini_ckpt]—they have limited impact for everyday LLM end-users. Model downloading is typically a one-time cost, so even if a model checkpoint is compressed by 50%, it only cuts the download time at most by half, presumably over the model’s entire lifecycle of deployment. Furthermore, checkpoints are usually stored on disk, where terabytes of capacity are easily available, making up a much looser constraint compared to GPU HBM (High Bandwidth Memory); one of the main resource constraints during inference.
@@
We argue that a lossless compression technique would be substantially more impactful if it could deliver efficiency gains during inference—particularly on GPU-based systems, which is the default setup for LLM serving. In this context, NeuZip [hao2024neuzip] is the only prior work we identify that supports GPU inference. NeuZip applies entropy encoding with layer-wise decompression to maintain a reduced memory footprint throughout serving. However, it is built on NVIDIA’s nvCOMP: “a high-speed data compression and decompression library optimized for NVIDIA GPUs”. [^6] https://developer.nvidia.com/nvcomp Unfortunately, nvCOMP is no longer open-source (only binary executables are available), which hinders future research. Moreover, we empirically find that nvCOMP’s inference throughput and latency are significantly worse than our proposed DFloat11 kernel, resulting in a pipeline that trades memory efficiency for substantial inference overhead (see Figure 7).
@@
Another work referencing NeuZip is Huff-LLM [yubeaton2025huffllm], which also aims to reduce memory costs while maintaining efficient inference. However, its contributions are specific to FPGA-like architectures and do not apply to GPUs. To the best of our knowledge, the DFloat data format we presented (and its respective kernel support in DFloat11) shall serve as the only GPU-inference-friendly data format with lossless compression benefits.
Efficient LLM Inference
LLMs are computationally intensive and resource-demanding, making the efficiency of LLM inference a key research focus [efficient_llm_survey]. FlashAttention [dao2022flashattention] accelerates exact attention computation on GPUs through kernel fusion, while NoMAD Attention [nomadattention] speeds up attention on CPUs using in-register lookups. Model compression is another effective strategy to reduce resource requirements for serving LLMs and diffusion models. Quantization methods such as GPTQ [frantar2022gptq], AWQ [lin2024awq], SmoothQuant [xiao2023smoothquant], LeanQuant [leanquant], CQ [cq], KVQuant [hooper2024kvquant], and KIVI [liu2024kivi] lower memory usage and enhance efficiency by compressing model weights, activations, or KV cache. Compression is also applied in fine-tuning: methods like LoRA [lora], QLoRA [dettmers2023qlora], and SketchTune [sketchtune] compress model weight deltas, whereas GaLore [galore] and SARA [sara] compress optimizer states during training. One additional line of work relevant to efficient LLM inference would be lossless efficient decoding, where paradigms such as speculative decoding [xia2022sd, leviathan2023sd, xia2024unlocking_sd_survey] and n-gram candidate decoding [fu2024lookahead, anonymous2025fafo] offer lossless generation quality with improved latency. DFloat11 mainly differs from these works in that it provides substantial savings in memory footprint while maintaining lossless generation quality, whereas most—if not all—lossless efficient decoding methods require memory consumption equal to or greater than that of the original model.